# Well-posedness of a first-order-in-time model for nonlinear acoustics with nonhomogeneous boundary conditions

### v2

@@ 2.3 Nonhomogeneous boundary conditions
Let Assumptions 2.1, 2.2 hold. Then, for $\tilde{r}>0$ small enough and any $f\in Y$, $g\in\tilde{W}$, $h\in X_{\mathfrak{T}}$ with $\tilde{h}(0)-g\in W$ such that
@@ 1. We wish to obtain $\tilde{u}$ as the unique solution of the linear problem
$\|\tilde{u}\|_{\tilde{X}}\leq C_{Z}(\|\partial_{t}\tilde{h}\|_{Y}+\|\tilde{\mathscr{A}}\|_{Y}+\|g\|_{\tilde{W}}+\|\tilde{h}(0)\|_{\tilde{W}})+c_{h}\|h\|_{X_{\mathfrak{T}}}\leq\tilde{C}(\|g\|_{\tilde{W}}+\|h\|_{X_{\mathfrak{T}}}),$
@@
$\mathcal{G}u^{0}:=(\partial_{t}u^{0}+\mathfrak{A}u^{0}+\mathfrak{B}[u^{0},u^{0}]+\tilde{\mathfrak{B}}[u^{0},\tilde{u}]+\tilde{\mathfrak{B}}[\tilde{u},u^{0}],u^{0}(0))-(f-\tilde{\mathfrak{B}}[\tilde{u},\tilde{u}],0).$
@@
$\mathcal{G}^{\prime}_{\tilde{{u_{*}}}}\cdot=(\partial_{t}\cdot+\mathfrak{A}\cdot+\mathfrak{B}[\cdot,\tilde{{u_{*}}}]+\mathfrak{B}[\tilde{{u_{*}}},\cdot]+\tilde{\mathfrak{B}}[\cdot,\tilde{u}]+\tilde{\mathfrak{B}}[\tilde{u},\cdot],\cdot(0))$
for each $\tilde{{u_{*}}}\in X$, since the only nonlinear term $\mathfrak{B}[u^{0},u^{0}]$ in $\mathcal{G}$ can be differentiated like the corresponding term in $\mathcal{F}$ in the proof of Theorem 2.2, see equation (2.2). The Lipschitz continuity of $\mathcal{G}^{\prime}$ with constant $K=2C_{\mathfrak{B}}C_{T}$ follows in the same way.
@@
An exception is that in the energy estimate of $\mathfrak{B}$, cf. (2.1.2), we now have to treat the term $\tilde{\mathfrak{B}}[u^{0},\tilde{u}]+\tilde{\mathfrak{B}}[\tilde{u},u^{0}]$. We can replace the $\|{u_{*}}\|_{X}$ terms by $\|\tilde{u}\|_{\tilde{X}}$ in (2.1.2) and the estimate can be carried out analogously in this case, since the assumptions (2.1(iii)), (2.1) are assumed to hold for the tilde spaces, respectively.
In the end, we obtain a unique $u^{0}\in X$ and constants $\tilde{r}_{G},\tilde{C_{G}}>0$ for which the a priori estimate $\|u^{0}\|_{X}\leq\tilde{C_{G}}\|f-\tilde{\mathfrak{B}}[\tilde{u},\tilde{u}]\|_{Y}$ is valid if $\tilde{u}\in B^{\tilde{X}}_{\tilde{r_{G}}}(0)$, which can be achieved due to (2.3) by assuming that $\|g\|_{\tilde{W}}+\|h\|_{X_{\mathfrak{T}}}$ is small enough.
@@
$\beta K\|-(\partial_{t}\cdot+\mathfrak{A}\cdot+\mathfrak{B}[\cdot,\tilde{u}]+\mathfrak{B}[\tilde{u},\cdot],\cdot(0))^{-1}(f-\tilde{\mathfrak{B}}[\tilde{u},\tilde{u}],0)\|_{X}\leq\tilde{C_{G}}^{2}K\|f-\tilde{\mathfrak{B}}[\tilde{u},\tilde{u}]\|_{Y}<\frac{1}{2}$
and $r^{-}<\tilde{r_{G}}$ can be ensured if $\tilde{r}$ is chosen small enough. This follows from smallness of $\|f\|_{Y}$ in combination with the boundedness of $\tilde{\mathfrak{B}}$ and the estimate (2.3).
It is left to show that $u=u^{0}+\tilde{u}$ is unique. If $\hat{u}$ is another solution to (2.3), then it can be decomposed into $\hat{u}=\hat{u}^{0}+\hat{\tilde{u}}$, where $\hat{\tilde{u}}$ satisfies (2.3) and $\hat{u}^{0}:=\hat{u}-\tilde{u}$ (2.3). Since these equations have unique solutions, it holds $u=\hat{u}$. ∎
@@ 3 Application to nonlinear acoustics
$\mathfrak{A}u=\mathcal{A}u+\mathscr{A}u=-\begin{pmatrix}\zeta\Delta p\\ \mu\Delta v\end{pmatrix}+\begin{pmatrix}\nabla\cdot v\\ \nabla p\end{pmatrix}$
@@
$\mathfrak{B}[u,z]=\begin{pmatrix}\alpha B_{1}[u,z]+\beta B_{2}[u,z]\\ \gamma B_{3}[u,z]+\delta B_{4}[u,z]\end{pmatrix}:=\begin{pmatrix}\alpha q\nabla\cdot v+\beta w\cdot\nabla p\\ \gamma q\nabla p+\delta(\nabla v)w\end{pmatrix}$,
@@ 3.1.1 Dirichlet Laplacian
The Dirichlet Laplacian $\Delta_{D}:{D}(-\Delta_{D})\subset L^{2}\to L^{2}$ can be obtained by a Friedrich’s extension, see [26, Theorem 10.19]. If we assume $\Omega$ to be $C^{2}$-regular, we have ${D}(-\Delta_{D})=H^{2}\cap H^{1}_{0}$ by elliptic regularity. Furthermore, $\Delta_{D}$ is self adjoint and has a spectral decomposition as required in assumption 2.1(ii) for $\mathcal{A}$ with $H=L^{2}$.
@@
${D}(-\Delta_{D}^{\frac{s}{2}})=\mathbb{H}^{s}:=\begin{cases}H^{s}\quad&0<s<\frac{1}{2},\\ H^{\frac{1}{2}}_{00}\quad&s=\frac{1}{2},\\ H^{s}_{0}\quad&\frac{1}{2}<s<1,\\ H^{s}\cap H^{1}_{0}\quad&1\leq s\leq 2,\end{cases}$
see [6, Theorem 3.1] in combination with [3, Proposition 4.1]. The space denoted by $H^{\frac{1}{2}}_{00}$ consists of all $u\in H^{\frac{1}{2}}$ such that $\int_{\Omega}d^{-1}u^{2}<\infty$, where $d(x):\Omega\to\mathbb{R}$ stands for the Euclidean distance from $x$ to $\partial\Omega$. We define $\mathbb{H}^{-s}:=(\mathbb{H}^{s})^{*}$.
Note that $\mathbb{H}^{s}\hookrightarrow H^{s}$ for $s\geq 0$ with embedding constant $C_{\mathbb{H}}>0$, which follows from [2, Proposition 2.2] and the fact that $H^{1}_{0}\hookrightarrow H^{1}$.
@@ 3.1.2 Neumann Laplacian
The classical Neumann Laplacian $\Delta_{N}:{D}(\Delta_{N})\subset L^{2}\to L^{2}$ can be obtained by a Friedrich’s extension, see [26, Theorem 10.20]. If $\Omega$ is assumed to be $C^{2}$-regular, we have
${D}(-\Delta_{N}):=\{u\in H^{2}:\mathcal{T}(\nabla u)\cdot n=0\},$
where $\mathcal{T}:H^{s}\to H^{s-\frac{1}{2}}(\partial\Omega)$ if $s>\frac{1}{2}$ denotes the classical trace operator, see [5, Chapter 5.5]. $\Delta_{N}$ is also self adjoint and has almost a spectral decomposition as required in assumption 2.1(ii) if $H=L^{2}$ with the exception that the first eigenvalue vanishes, i.e., $\lambda_{1}=0$. This can be overcome by considering $\Delta_{N}:{D}(-\Delta_{N})_{\diagup{C}}\subset L^{2}_{\diagup{C}}\to L^{2}_{\diagup{C}}$, i.e., by factoring out constant functions. We write for this operator $\Delta_{\tilde{N}}$ with ${D}(-\Delta_{\tilde{N}})={D}(-\Delta_{N})_{\diagup{C}}$. Then, this Neumann Laplacian $\Delta_{\tilde{N}}$ does not have a vanishing eigenvalue, since the eigenspace corresponding to $\lambda_{1}$ for $\Delta_{N}$ equals $C$.
@@
${D}(-\Delta_{\tilde{N}}^{\frac{s}{2}})=\mathbb{G}^{s}:=\begin{cases}\{u\in H^{s}:\int_{\Omega}u=0\}\quad&s<\frac{3}{2},\\ \{u\in H^{\frac{3}{2}}:\int_{\Omega}u=0,\ |\nabla u|\in H^{\frac{1}{2}}_{00}\}\quad&s=\frac{3}{2},\\ \{u\in H^{s}:\int_{\Omega}u=0,\ \mathcal{T}(\nabla u)\cdot n=0\}\quad&s>\frac{3}{2},\end{cases}$
see [2, Proposition 2.4] for $s<1$ and [14, Remark 2.10] for $s\geq 1$. Note that the condition $\int_{\Omega}u=0$ is equivalent to factoring out constant functions, and we define $\mathbb{G}^{-s}:=(\mathbb{G}^{s})^{*}$.
@@ 3.1.3 Hodge/Lions Laplacian
For Hodge/Lions boundary conditions, we only cover the case $d=3$ for simplicity, so that the vector cross product $\times$ is a map $\times:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$. We consider a vector valued Laplacian acting on functions with $d=3$ components fulfilling the Hodge/Lions boundary conditions as in [20, 21], denoted by $\Delta_{M}$ in the following. This operator can be constructed via a self adjoint Friedrich’s extension with domain
${D}(-\Delta_{M})=\{v\in H^{1}_{\perp}:\nabla\cdot v\in H^{1},\ \nabla\times(\nabla\times v)\in(L^{2})^{d},\ (\nabla\times v)\times n=0\text{ on }\partial\Omega\}$
@@
in case of a Lipschitz-regular domain $\Omega$, see [21, Section 4]. When $\Omega$ is assumed to be more regular, one can obtain more regular function spaces. This is shown in the next Lemma, justifying our notation $H^{1}_{\perp}$. Note that this construction of $\Delta_{M}$ and the representation of $D(-\Delta_{M})$ is linked to the vector calculus formula $\Delta v=\nabla\nabla\cdot v-\nabla\times\nabla\times v$.
@@
$\{v\in(H^{1})^{d}:\nabla\cdot v\in H^{1},\ \nabla\times v\in(H^{1})^{d},\ v\cdot n=c\text{ on }\partial\Omega\}\hookrightarrow(H^{2})^{d}$
@@
The embeddings (3.1), (3.1) can be deduced from [1] using [1, Theorems 2.9, 2.12, Remark 2.14 and Corollary 2.15 with $m=2$].
To prove ${D}(-\Delta_{M})\hookrightarrow(H^{2})^{d}$ note that for $w:=\nabla\times v$ with $v\in{D}(-\Delta_{M})$ we have $\nabla\cdot w=\nabla\cdot\nabla\times v=0\in L^{2}$ (in a distributional sense), $\nabla\times w=\nabla\times\nabla\times v\in(L^{2})^{d}$ and $w\times n=\nabla\times v\times n=0$ on $\partial\Omega$. Therefore, applying (3.1) gives us $w\in(H^{1})^{d}$, that is $\nabla\times v\in(H^{1})^{d}$. Thus, using (3.1) we obtain ${D}(-\Delta_{M})\hookrightarrow(H^{2})^{d}$. ∎
@@
Finally, the following Lemma is a well-posedness result for $\Delta_{M}$ following from [20, Theorem 1.2].
@@
The statement is a special case of [20, Theorem 1.2]. Within the notation of [20], we set $p=2$ and choose $\mathcal{M}=\mathbb{R}^{3}$ with Euclidean metric as manifold. Then $\Omega$ is a Lipschitz domain of $\mathbb{R}^{3}$ and in the case of $p=2$ and $C^{1}$-regular $\Omega$ the valid choices of $s$ in [20, Theorem 1.2] according to [20, (1.1)] are characterized by
$\frac{1}{2}<s<\frac{1}{2},\quad 0<\frac{1}{2}-\frac{s}{3}<1.$
The estimate (3.2) follows from the fact that according to [20, Theorem 1.2] the boundary value problem (3.2) is Fredholm solvable of index zero, which in particular implies that the solution operator $(\chi,\xi,\nu)\mapsto v$ corresponding to (3.2) is bounded. ∎

### v3

@@ 2.3 Nonhomogeneous boundary conditions
Let $T>0$ and Assumptions 2.1, 2.2 hold. Let $f\in Y$, $g\in\tilde{W}$, $h\in X_{\mathfrak{T}}$ with $\tilde{h}(0)-g\in W$. Then there exists a $\tilde{r}>0$ such that if
@@ 1. We wish to obtain $\tilde{u}$ as the unique solution of the linear problem
$\|\tilde{u}\|_{\tilde{X}}\leq C_{Z}(\|\partial_{t}\tilde{h}\|_{Y}+\|\tilde{\mathscr{A}}\|_{Y}+\|g\|_{\tilde{W}}+\|\tilde{h}(0)\|_{\tilde{W}})+c_{h}\|h\|_{X_{\mathfrak{T}}}\leq\tilde{C}(\|g\|_{\tilde{W}}+\|h\|_{X_{\mathfrak{T}}}),$ (36)
@@
$\mathcal{G}u^{0}:=(\partial_{t}u^{0}+\mathfrak{A}u^{0}+\mathfrak{B}[u^{0},u^{0}]+\tilde{\mathfrak{B}}[u^{0},\tilde{u}]+\tilde{\mathfrak{B}}[\tilde{u},u^{0}],u^{0}(0))-(f-\tilde{\mathfrak{B}}[\tilde{u},\tilde{u}],0).$ (37)
@@
$\mathcal{G}^{\prime}_{\tilde{{u_{*}}}}\cdot=(\partial_{t}\cdot+\mathfrak{A}\cdot+\mathfrak{B}[\cdot,\tilde{{u_{*}}}]+\mathfrak{B}[\tilde{{u_{*}}},\cdot]+\tilde{\mathfrak{B}}[\cdot,\tilde{u}]+\tilde{\mathfrak{B}}[\tilde{u},\cdot],\cdot(0))$ (38)
for each $\tilde{{u_{*}}}\in X$, since the only nonlinear term $\mathfrak{B}[u^{0},u^{0}]$ in $\mathcal{G}$ can be differentiated like the corresponding term in $\mathcal{F}$ in the proof of Theorem 2.2, see equation (32). The Lipschitz continuity of $\mathcal{G}^{\prime}$ with constant $K=2C_{\mathfrak{B}}C_{T}$ follows in the same way.
@@
An exception is that in the energy estimate of $\mathfrak{B}$, cf. (2.1.2), we now have to treat the term $\tilde{\mathfrak{B}}[u^{0},\tilde{u}]+\tilde{\mathfrak{B}}[\tilde{u},u^{0}]$. We can replace the $\|{u_{*}}\|_{X}$ terms by $\|\tilde{u}\|_{\tilde{X}}$ in (2.1.2) and the estimate can be carried out analogously in this case, since the assumptions (11), (16) are assumed to hold for the tilde spaces, respectively.
In the end, we obtain a unique $u^{0}\in X$ and constants $\tilde{r}_{G},\tilde{C_{G}}>0$ for which the a priori estimate $\|u^{0}\|_{X}\leq\tilde{C_{G}}\|f-\tilde{\mathfrak{B}}[\tilde{u},\tilde{u}]\|_{Y}$ is valid if $\tilde{u}\in B^{\tilde{X}}_{\tilde{r_{G}}}(0)$, which can be achieved due to (36) by assuming that $\|g\|_{\tilde{W}}+\|h\|_{X_{\mathfrak{T}}}$ is small enough.
@@
$\beta K\|-(\partial_{t}\cdot+\mathfrak{A}\cdot+\mathfrak{B}[\cdot,\tilde{u}]+\mathfrak{B}[\tilde{u},\cdot],\cdot(0))^{-1}(f-\tilde{\mathfrak{B}}[\tilde{u},\tilde{u}],0)\|_{X}\leq\tilde{C_{G}}^{2}K\|f-\tilde{\mathfrak{B}}[\tilde{u},\tilde{u}]\|_{Y}<\frac{1}{2}$ (39)
and $r^{-}<\tilde{r_{G}}$ can be ensured if $\tilde{r}$ is chosen small enough. This follows from smallness of $\|f\|_{Y}$ in combination with the boundedness of $\tilde{\mathfrak{B}}$ and the estimate (36).
It is left to show that $u=u^{0}+\tilde{u}$ is unique. If $\hat{u}$ is another solution to (2.3), then it can be decomposed into $\hat{u}=\hat{u}^{0}+\hat{\tilde{u}}$, where $\hat{\tilde{u}}$ satisfies (2.3) and $\hat{u}^{0}:=\hat{u}-\tilde{u}$ (2.3). Since these equations have unique solutions, it holds $u=\hat{u}$.
@@ 3 Application to nonlinear acoustics
$\mathfrak{A}u=\mathcal{A}u+\mathscr{A}u=-\begin{pmatrix}\zeta\Delta p\\ \mu\Delta v\end{pmatrix}+\begin{pmatrix}\nabla\cdot v\\ \nabla p\end{pmatrix}$ (40)
@@
$\mathfrak{B}[u,z]=\begin{pmatrix}\alpha B_{1}[u,z]+\beta B_{2}[u,z]\\ \gamma B_{3}[u,z]+\delta B_{4}[u,z]\end{pmatrix}:=\begin{pmatrix}\alpha q\nabla\cdot v+\beta w\cdot\nabla p\\ \gamma q\nabla p+\delta(\nabla v)w\end{pmatrix}$, (41)
@@ 3.1.1 Dirichlet Laplacian
The Dirichlet Laplacian $\Delta_{D}:{D}(-\Delta_{D})\subset L^{2}\to L^{2}$ can be obtained by a Friedrich’s extension, see [schmudgen2012, Theorem 10.19]. If we assume $\Omega$ to be $C^{2}$-regular, we have ${D}(-\Delta_{D})=H^{2}\cap H^{1}_{0}$ by elliptic regularity. Furthermore, $\Delta_{D}$ is self adjoint and has a spectral decomposition as required in assumption 2.1(ii) for $\mathcal{A}$ with $H=L^{2}$.
@@
${D}(-\Delta_{D}^{\frac{s}{2}})=\mathbb{H}^{s}:=\begin{cases}H^{s}\quad&0<s<\frac{1}{2},\\ H^{\frac{1}{2}}_{00}\quad&s=\frac{1}{2},\\ H^{s}_{0}\quad&\frac{1}{2}<s<1,\\ H^{s}\cap H^{1}_{0}\quad&1\leq s\leq 2,\end{cases}$ (43)
see [fefferman2022, Theorem 3.1] in combination with [bonito2013, Proposition 4.1]. The space denoted by $H^{\frac{1}{2}}_{00}$ consists of all $u\in H^{\frac{1}{2}}$ such that $\int_{\Omega}d^{-1}u^{2}<\infty$, where $d(x):\Omega\to\mathbb{R}$ stands for the Euclidean distance from $x$ to $\partial\Omega$. We define $\mathbb{H}^{-s}:=(\mathbb{H}^{s})^{*}$.
Note that $\mathbb{H}^{s}\hookrightarrow H^{s}$ for $s\geq 0$ with embedding constant $C_{\mathbb{H}}>0$, which follows from [antil2017, Proposition 2.2] and the fact that $H^{1}_{0}\hookrightarrow H^{1}$.
@@ 3.1.2 Neumann Laplacian
The classical Neumann Laplacian $\Delta_{N}:{D}(\Delta_{N})\subset L^{2}\to L^{2}$ can be obtained by a Friedrich’s extension, see [schmudgen2012, Theorem 10.20]. If $\Omega$ is assumed to be $C^{2}$-regular, we have
${D}(-\Delta_{N}):=\{u\in H^{2}:\mathcal{T}(\nabla u)\cdot n=0\},$ (44)
where $\mathcal{T}:H^{s}\to H^{s-\frac{1}{2}}(\partial\Omega)$ if $s>\frac{1}{2}$ denotes the classical trace operator, see [Evans:2010, Chapter 5.5]. $\Delta_{N}$ is also self adjoint and has almost a spectral decomposition as required in assumption 2.1(ii) if $H=L^{2}$ with the exception that the first eigenvalue vanishes, i.e., $\lambda_{1}=0$. This can be overcome by considering $\Delta_{N}:{D}(-\Delta_{N})_{\diagup{C}}\subset L^{2}_{\diagup{C}}\to L^{2}_{\diagup{C}}$, i.e., by factoring out constant functions. We write for this operator $\Delta_{\tilde{N}}$ with ${D}(-\Delta_{\tilde{N}})={D}(-\Delta_{N})_{\diagup{C}}$. Then, this Neumann Laplacian $\Delta_{\tilde{N}}$ does not have a vanishing eigenvalue, since the eigenspace corresponding to $\lambda_{1}$ for $\Delta_{N}$ equals $C$.
@@
${D}(-\Delta_{\tilde{N}}^{\frac{s}{2}})=\mathbb{G}^{s}:=\begin{cases}\{u\in H^{s}:\int_{\Omega}u=0\}\quad&s<\frac{3}{2},\\ \{u\in H^{\frac{3}{2}}:\int_{\Omega}u=0,\ |\nabla u|\in H^{\frac{1}{2}}_{00}\}\quad&s=\frac{3}{2},\\ \{u\in H^{s}:\int_{\Omega}u=0,\ \mathcal{T}(\nabla u)\cdot n=0\}\quad&s>\frac{3}{2},\end{cases}$ (45)
see [antil2017, Proposition 2.4] for $s<1$ and [kim2020, Remark 2.10] for $s\geq 1$. Note that the condition $\int_{\Omega}u=0$ is equivalent to factoring out constant functions, and we define $\mathbb{G}^{-s}:=(\mathbb{G}^{s})^{*}$.
@@ 3.1.3 Hodge/Lions Laplacian
For Hodge/Lions boundary conditions, we only cover the case $d=3$ for simplicity, so that the vector cross product $\times$ is a map $\times:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$. We consider a vector valued Laplacian acting on functions with $d=3$ components fulfilling the Hodge/Lions boundary conditions as in [mitrea2004, monniaux2013], denoted by $\Delta_{M}$ in the following. This operator can be constructed via a self adjoint Friedrich’s extension with domain
${D}(-\Delta_{M})=\{v\in H^{1}_{\perp}:\nabla\cdot v\in H^{1},\ \nabla\times(\nabla\times v)\in(L^{2})^{d},\ (\nabla\times v)\times n=0\text{ on }\partial\Omega\}$ (46)
@@
in case of a Lipschitz-regular domain $\Omega$, see [monniaux2013, Section 4]. When $\Omega$ is assumed to be more regular, one can obtain more regular function spaces. This is shown in the next Lemma, justifying our notation $H^{1}_{\perp}$. Note that this construction of $\Delta_{M}$ and the representation of $D(-\Delta_{M})$ is linked to the vector calculus formula $\Delta v=\nabla\nabla\cdot v-\nabla\times\nabla\times v$.
@@
$\{v\in(H^{1})^{d}:\nabla\cdot v\in H^{1},\ \nabla\times v\in(H^{1})^{d},\ v\cdot n=c\text{ on }\partial\Omega\}\hookrightarrow(H^{2})^{d}$ (47)
@@
The embeddings (3.1), (47) can be deduced from [amrouche1998] using [amrouche1998, Theorems 2.9, 2.12, Remark 2.14 and Corollary 2.15 with $m=2$].
To prove ${D}(-\Delta_{M})\hookrightarrow(H^{2})^{d}$ note that for $w:=\nabla\times v$ with $v\in{D}(-\Delta_{M})$ we have $\nabla\cdot w=\nabla\cdot\nabla\times v=0\in L^{2}$ (in a distributional sense), $\nabla\times w=\nabla\times\nabla\times v\in(L^{2})^{d}$ and $w\times n=\nabla\times v\times n=0$ on $\partial\Omega$. Therefore, applying (3.1) gives us $w\in(H^{1})^{d}$, that is $\nabla\times v\in(H^{1})^{d}$. Thus, using (47) we obtain ${D}(-\Delta_{M})\hookrightarrow(H^{2})^{d}$. ∎
@@
Finally, the following Lemma is a well-posedness result for $\Delta_{M}$ following from [mitrea2004, Theorem 1.2].
@@
The statement is a special case of [mitrea2004, Theorem 1.2]. Within the notation of [mitrea2004], we set $p=2$ and choose $\mathcal{M}=\mathbb{R}^{3}$ with Euclidean metric as manifold. Then $\Omega$ is a Lipschitz domain of $\mathbb{R}^{3}$ and in the case of $p=2$ and $C^{1}$-regular $\Omega$ the valid choices of $s$ in [mitrea2004, Theorem 1.2] according to [mitrea2004, (1.1)] are characterized by
$\frac{1}{2}<s<\frac{1}{2},\quad 0<\frac{1}{2}-\frac{s}{3}<1.$ (48)
The estimate (3.2) follows from the fact that according to [mitrea2004, Theorem 1.2] the boundary value problem (3.2) is Fredholm solvable of index zero, which in particular implies that the solution operator $(\chi,\xi,\nu)\mapsto v$ corresponding to (3.2) is bounded.